OpenStudy (anonymous):

The circumference of a circle is 15 in. The diameter of the circle is _____ in

OpenStudy (ganpat):

circumference is measure of outer part.. So C = 2 * pi * r 15 = 2 * 22/7 * r r = (7 *15)/ (2 * 22) r = 2.38 in...